OpenStudy (anonymous):

find the exact value of: arctan(-square root of 3) answer in radians in terms of pi

OpenStudy (jamesj):

well .... cos(pi/6) = 1/2 and sin(pi/6) = sqrt(3)/ 2 => tan(pi/6) = sqrt(3) and tan(-pi/6) = -sqrt(3) because sin(-pi/6) = - sqrt(3)/2 and cos(-pi/6) = 1/2

OpenStudy (anonymous):

look at unit circle on last page of cheat sheet. if a point on the circle where the ratio of the second coordinate to the first is \[-\sqrt{3}\] and then make sure to answer with an angle (number) in the interval \[[-\frac{\pi}{2},\frac{\pi}{2}]\] in other words on the right half of the circle